If a hog cause 15cent per pound and its a 150 pound how much would it cost?
Answer by jim_thompson5910(35256)

(15 cents/1 pound)*(150 pounds) = 15*150 = 2,250 cents

2,250 cents = 2,250*(1 dollar/100 cents) = 2,250/100 = 22.50 dollars

It costs $22.50